OpenStudy (sarahc):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 33 feet and the child ride has a radius of 22 feet, what is the ratio between the volumes of each ride? 27:8 9:6 9:4 3:2

OpenStudy (hba):

radius ratio=33/22 volume ratio=(radius ratio)^3

OpenStudy (sarahc):

What would be the answer? :)

OpenStudy (hba):

solve it urself n tell me the answer n ill tell u wheter ur right or wrong

OpenStudy (sarahc):

9:4 ?

OpenStudy (hba):

did u solve it

OpenStudy (sarahc):

errr was I wrong?

OpenStudy (hba):

27:8 is the answer i guess

OpenStudy (sarahc):

I just got that....lol 9/4=3/2 then 3^3/2^3 = 27/8

OpenStudy (sarahc):

i mean 27:8

OpenStudy (hba):

good

OpenStudy (sarahc):

thanks

OpenStudy (hba):

ur welcome